OpenStudy (anonymous):

if x+x=2 and how x the power of 3 plus x the power of 3 ? a.2 b.4 c.6 d.8

OpenStudy (anonymous):

how 6+6+6+6+6+6x0 ?

OpenStudy (anonymous):

a. 2 x+x=2 plug in 1 for x 1+1=2 (1)^3 + (1)^3 = 2

OpenStudy (anonymous):

6+6+6+6+6+6x0 6+6+6+6+6+0 6+6+6+6+6 5(6) 30